OpenStudy (anonymous):

1. Which of the following is an example of an exponential function? A. f(x) = 4 B. f(x) = 4x C. f(x)= 4/x D. f(x) = 6x

OpenStudy (solomonzelman):

Which one do you think?

OpenStudy (anonymous):

umm maybe C

OpenStudy (solomonzelman):

Yes, you can exclude A which is just a horizontal line y=4, Exclude B and C that are leniar lines with slopes of 4 and 6 respectively. Also, looking at C \(\LARGE\color{blue}{ \bf \frac{4}{1},~~ }\)\(\LARGE\color{blue}{ \bf \frac{4}{2},~~ }\)\(\LARGE\color{blue}{ \bf \frac{4}{3},~~ }\)\(\LARGE\color{blue}{ \bf \frac{4}{4},~~ }\) \(\LARGE\color{blue}{ \bf \frac{4}{5}....~~ }\)

OpenStudy (solomonzelman):

you can see the exponential decay.